OpenStudy (anonymous):

Out of a swarm of bees one fifth part settled on a Kadamba blossom; one third on a Silihindra flower; three times the difference of those numbers flew to the bloom of a Kutaja. One bee, which remained, hovered about in the air. Tell me, charming woman, the number of bees??? Hint: Call the number of bees and convert the story into an equation.

OpenStudy (anonymous):

(1/3)x = KB (1/5)x = SF KB - SF = y 3y = K x = total number of bees x - 1 = KB + SF + K x - K - KB - SF = 1 x - 3([1/3]x - [1/5]x) - 1/3x - 1/5x = 1 x - x + 3/5x - 1/3x - 1/5x 2/5x - 1/3x = 1 6/15 - 5/15 = 1 1/15 = 1 x = 15